How do I Solve by completing the square:
3x^2 + x + 1 = 0

Answer by stanbon(75887)

Solve by completing the square:
3x^2 + x + 1 = 0
------
3x^2 + x = -1
----
3(x^2 + x/3 + (1/9) = -1 + 3(1/9)
------
3(x+1/3)^2 = 2
---
(x+1/3)^2 = 2/3
------
x+(1/3) = sqrt(2/3) or x+(1/3) = -sqrt(2/3)
----
x = [-1+sqrt(6)]/3 or x = [-1-sqrt(6)]/3
-------------
Cheers,
Stan H.
------------